OpenStudy (anonymous):

Find the inverse of the function. f(x) = 7x + 8 @satellite73 @jim_thompson5910 @zepdrix @dan815 @iambatman @thesmartone @perl @kainui

OpenStudy (anonymous):

can you solve \[7x+8=20\]?

jimthompson5910 (jim_thompson5910):

Basic Idea: 1) replace f(x) with y 2) swap x and y 3) solve for y to get the inverse ------------------------------------------------------------------------ \[f(x) = 7x + 8\] \[y = 7x + 8\] \[x = 7y + 8\] \[x-8 = 7y\] what's the next step?

OpenStudy (anonymous):

i have no clue what to do next @jim_thompson5910

jimthompson5910 (jim_thompson5910):

7y means 7 times y

jimthompson5910 (jim_thompson5910):

what is the complete opposite of multiplication?

OpenStudy (anonymous):

division @jim_thompson5910

jimthompson5910 (jim_thompson5910):

so you divide both sides by 7

jimthompson5910 (jim_thompson5910):

Basic Idea: 1) replace f(x) with y 2) swap x and y 3) solve for y to get the inverse ------------------------------------------------------------------------ \[f(x) = 7x + 8\] \[y = 7x + 8\] \[x = 7y + 8\] \[x-8 = 7y\] \[\frac{x-8}{7} = y\] \[y = \frac{x-8}{7}\] \[f^{-1}(x) = \frac{x-8}{7}\]

OpenStudy (anonymous):

ok so its y=8-x/7 @jim_thompson5910

OpenStudy (anonymous):

is that the final answer @jim_thompson5910

jimthompson5910 (jim_thompson5910):

yes what I got is the correct inverse

jimthompson5910 (jim_thompson5910):

you can also state it as \[f^{-1}(x) = \frac{1}{7}x - \frac{8}{7}\]

OpenStudy (anonymous):

thank u can you help me with another one @jim_thompson5910

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Find the inverse of the function. f(x) = 8x3 - 5 @jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\Large f(x) = 8x^3 - 5\] right?

OpenStudy (anonymous):

yes @jim_thompson5910

jimthompson5910 (jim_thompson5910):

same idea as last time 1) replace f(x) with y 2) swap x and y 3) solve for y to get the inverse show me how far you get when you follow these steps

OpenStudy (anonymous):

y = 8x^3-5 8x^3=y-5 i dont know how to continue the rest @jim_thompson5910

jimthompson5910 (jim_thompson5910):

you undo the -5 by adding 5 to both sides

jimthompson5910 (jim_thompson5910):

try again

OpenStudy (anonymous):

im so confused can you show me @jim_thompson5910

jimthompson5910 (jim_thompson5910):